OpenStudy (anonymous):

Which of the following is NOT \[\sqrt{8(\cos \frac{ 11\pi }{ 6 } + i \sin \frac{ 11\pi }{ 6 })}\] a) \[2^{3/2}(\cos \frac{ 11\pi }{ 12 } + \sin \frac{ 11\pi }{ 12 })\] b) \[2^{3/2}(-\cos \frac{ \pi }{ 12 } +i \sin \frac{ \pi }{ 12 })\] c) \[8^{1/2}(\cos \frac{ 23\pi }{ 12 } +i \sin \frac{ 23\pi }{ 12 })\] d) \[8^{1/2}(\cos \frac{ \pi }{ 12 } -i \sin \frac{ \pi }{ 12 })\] e) \[2^{3/2}(\cos \frac{23 \pi }{ 12 } -i \sin \frac{ 23\pi }{ 12 })\]

OpenStudy (anonymous):

r u doin a test?

OpenStudy (anonymous):

Exam prep.

OpenStudy (anonymous):

We can't help u sry...

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Have a great day sir

OpenStudy (anonymous):

Thank you, you 2!

OpenStudy (helder_edwin):

use DeMoivre's theorem \[ \large [\rho(\cos\phi+i\sin\phi)]^n=\rho^n(\cos(n\phi)+i\sin(n\phi)) \]

OpenStudy (anonymous):

But what is the n value in this case? since it is under a square root?

OpenStudy (helder_edwin):

n=1/2

OpenStudy (anonymous):

May I ask how you find this?

OpenStudy (helder_edwin):

\[ \large \sqrt{8(\cos\frac{11\pi}{6}+i\sin\frac{11\pi}{6})}= \left(8(\cos\frac{11\pi}{6}+i\sin\frac{11\pi}{6})\right)^{1/2} \] \[ \large =8^{1/2}(\cos\frac{1}{2}\frac{11\pi}{6}+i\sin\frac{1}{2}\frac{11\pi}{6}) \]

OpenStudy (anonymous):

Thus option e is the correct one. (The one that is not equal to the above). Since it doesn't give -2?

OpenStudy (helder_edwin):

\[ \large =(2^3)^{1/2}(\cos\frac{11\pi}{12}+i\sin\frac{11\pi}{12}) \]

OpenStudy (anonymous):

@ helder, I really don't catch on as to what you are trying to tell me now. Don't you agree with my answer?

OpenStudy (helder_edwin):

sorry. this site aparently crashed. and i had to go to work. the correct answer is A.